Make the variable in the bracket the subject of the given formula.
1. V=(1/3)πr^2h {r}
1
Expert's answer
2012-09-11T10:27:33-0400
V=(1/3)πr^2h
whence r^2=3V/(πh)
r=sqrt(3V/(πh)) (assuming that r>=0)